# L-PR: Exploiting LiDAR Fiducial Marker for Unordered Low Overlap Multiview Point Cloud Registration

### v2

@@ III-C First-level Graph
Specifically, our objective is to infer the relative poses with the highest possible accuracy using only the necessary low-dimensional information through a simple process. Thus, we construct the first-level graph as a weighted graph. The construction of this graph is as follows. When processing the point clouds with the proposed adaptive threshold method one by one, if a marker is detected in a point cloud, we add the corresponding scan node and marker node to the graph along with a weighted edge. The edge weight is the pose estimation point-to-point error ${e}_{pp}$:
@@
Eq. (5) indicates the substitution of the result given by SVD back into the right side of Eq. (2). ${e}_{pp}\in\mathbb{R}^{+}$ is employed as the metric to evaluate the quality of pose estimation of the marker in the corresponding point cloud. We define the first scan as the anchor scan. The local coordinate system of the anchor scan is set as the global coordinate system. Namely, $\{F\}_{1}=\{G\}$. We only consider the relative poses between the anchor scan and each non-anchor scan. Although there could be multiple paths from the given scan to the anchor scan (see Fig. 4), we have already saved the information on pose estimation quality by constructing the first-level graph as a weighted graph.
Figure 4: An illustration of the first-level graph. In this simple case, as shown on the right side, four alternative paths exist from $f_{2}$ to the anchor scan ($f_{1}$). We employ Dijkstra’s algorithm [30] to find the shortest path.
Therefore, we adopt Dijkstra’s algorithm [30] to obtain the shortest path. Then, we compute the relative pose along the shortest path iteratively using Eq. (4). Given that the relative pose computation is achieved with the lowest accumulation of $e_{pp}$, the estimation result achieves the highest possible accuracy. Moreover, the search for the optimal path to propagate poses is based merely on the one-dimensional $e_{pp}$, without incorporating any 6-DOF poses or 3D locations.
@@
III-D Second-level Graph
In this section, we construct a factor graph to optimize the variables, ultimately addressing the MAP inference problem introduced in Eq. (3). Suppose that the $j$-th marker is detected in the $i$-th scan, the following nodes are added to the second graph. The variable nodes include $\mathbf{T}^{j}$ (6-DOF pose of the $j$-th marker w.r.t. $\{G\}$), $\mathbf{T}_{i}$ (6-DOF pose of the $i$-th scan w.r.t. $\{G\}$), and $\{\mathbf{p}^{j,1},\mathbf{p}^{j,2},\mathbf{p}^{j,3},\mathbf{p}^{j,4}\}$ (3D coordinates of the corners of the $j$-th marker w.r.t. $\{G\}$). The factor nodes include $\mathbf{T}^{j}_{i}$ (6-DOF pose of the $j$-th marker w.r.t. $\{F\}_{i}$), $\{^{j}\mathbf{p}^{j,1},\ ^{j}\mathbf{p}^{j,2},\ ^{j}\mathbf{p}^{j,3},\ ^{j}\mathbf{p}^{j,4}\}$ (3D coordinates of the coroners of the $j$-th marker w.r.t. $\{M\}_{j}$), and $\{_{i}\mathbf{p}^{j,1},\ _{i}\mathbf{p}^{j,2},$ ${}_{i}\mathbf{p}^{j,3},$ ${}_{i}\mathbf{p}^{j,4}\}$ (3D coordinates of the corners of the $j$-th marker w.r.t. $\{F\}_{i}$). Note that by adding variables representing the 3D coordinates of the corners, we can also optimize the fiducial localization results at this stage. For clarity, the first stage in Fig. 5 shows the added nodes and edges when a marker is detected in a scan. The graph becomes the one depicted in the second stage of Fig. 5 after integrating all the marker detection results.
Figure 5: The procedures for formulating the second graph. The variable nodes are represented by circles and the factor nodes are represented by squares.
Subsequently, we add a prior factor that connects to the first (anchor) point cloud pose node since we set $\{F\}_{1}$ as $\{G\}$. Finally, we add factor nodes representing the relative poses between the anchor scan and each non-anchor scan. Up to this point, the factor graph is completed, as shown in stage 3 of Fig. 5. Following [31, 32], this indicates that the factorization of the posterior probability $P(\Theta\mid\mathcal{Z})$ has been defined:
@@ III-D Second-level Graph
where $P^{(k)}$ are the factors in the second graph and we follow the standard pipeline [31] to model them as Gaussians:
@@
where $h_{k}\left(\Theta\right)$ is the measurement function and $z_{k}$ is a measurement. $\|e\|_{\Sigma}^{2}\triangleq e^{T}\Sigma^{-1}e$ denotes the squared Mahalanobis distance with $\Sigma$ being the covariance matrix. Following [23], if $z_{k}\in\mathbb{R}^{3\times 1}$ is a 3D position, $\ominus$ refers to straight subtraction for elements. If $z_{k}\in SE(3)$ is a 6-DOF pose, $\ominus$ generates 6-dimensional Lie algebra coordinates:
@@
where for a $\mathbf{T}\in SE(3)$, $\mathrm{Rot}(\mathbf{T})$ denotes the rotation matrix $\mathbf{R}\in SO(3)$ and $\mathrm{Trans}(\mathbf{T})$ denotes translation vector $\mathbf{t\in\mathbb{R}^{3\times 1}}$. $\log(\cdot)$ represents the matrix logarithm. $\vee$ is the vee map operator that finds the unique vector $\xi\in\mathbb{R}^{3\times 1}$ (Lie algebra coordinates) corresponding to a given skew-symmetric matrix $\log(\mathbf{R})\in\mathbb{R}^{3\times 3}$ [29, 23]. With the Gaussian modeling, the objective function in Eq. (6) is transformed into a least square problem by applying the negative logarithm:
@@
We utilize the Levenberg-Marquardt algorithm [32] to solve this problem. The acquisition of initial values is introduced in Section III-C. The noise covariance matrices are determined by the quantitative experiments conducted in [19].
IV Livox-3DMatch Dataset
A common approach [1, 2] to evaluating a learning-based point cloud registration model is to train it on 3DMatch [6] and test it on various benchmarks, including 3DMatch [6], ETH [21], and ScanNet [22]. However, the 3DMatch benchmark is mainly constructed from RGB-D camera captures of indoor scenes [6]. In this work, we collect a new dataset named Livox-3DMatch to enrich the training data for learning-based methods. The enrichment contains two key components. Firstly, the sensor we adopt is a Livox MID-40 LiDAR, which has different sampling patterns compared to the RGB-D camera (see Fig. 6(a) and (b)). Thus, it adds point clouds with new features to the training data. Secondly, we selectively sampled some scenes that are absent or rare in 3DMatch, thereby enriching the scenes in the training data. For example, the valid depth range of an RGB-D camera is usually less than ten meters, making it unsuitable for sampling outdoor scenes. In contrast, we collect some outdoor scenes (see Fig. 6(c)) for Livox-3DMatch, considering that a LiDAR can sample objects a few hundred meters away. Moreover, we collect some challenging cases (see Fig. 6(d)) where the overlapping regions are mainly planes, which are rare in 3DMatch [6]. A more detailed introduction to the scenes in Livox-3DMatch is given in Sections V-B and V-C. In Section V-D, we demonstrate that the proposed Livox-3DMatch can boost the performance of the SOTA learning-based method [1] on various benchmarks [6, 21, 22].
@@ IV Livox-3DMatch Dataset
V Experimental Results
In this section, we report the experimental results of seven tests to demonstrate the effectiveness and applications of the proposed method from different perspectives.
V-A Evaluation of the Adaptive Marker Detection Method
Data. We place the Livox MID-40 LiDAR and 69.2 cm $\times$ 69.2 cm ArUco marker(s) in three different scenes, as shown in Fig. 7. The three scenes are: between groups of buildings, in open outdoor areas, and in large indoor parking lots. In each scene, we collect point clouds and test the proposed adaptive threshold method at different relative positions. Results and Analysis. Table I presents the results, where x and y indicate the relative position of the LiDAR in the marker (ID 4) coordinate system. Specifically, in scene Fig. 7(c), there are two markers (ID 4 and ID 1), so the column for $\lambda^{*}$ contains two values. As seen in the table, $\lambda^{*}$ varies significantly as the relative position changes, demonstrating the necessity of the proposed adaptive marker detection method. In addition, the results shown in Fig. 7(c) also illustrate the necessity of the memory queue design, considering that a single $\lambda^{*}$ might not be applicable to all markers in the same scene.
@@ V-A Evaluation of the Adaptive Marker Detection Method
TABLE I: Demonstration of the necessity of the proposed adaptive marker detection algorithm
@@
V-B Evaluation of Point Cloud Registration Accuracy
Data. Given that the existing point cloud registration benchmarks [6, 21, 22] lack fiducial markers in the scenes, we construct a new dataset, as shown in Fig. 8, with the Livox MID-40. As illustrated in the caption of Fig. 8, our dataset cover various indoor and outdoor scenes. Indoors, multiple 16.4 cm × 16.4 cm AprilTags [10] are positioned in the environment. Outdoors, multiple 69.2 cm × 69.2 cm ArUcos [11] are placed in the environment. Note that since the LiDAR fiducial markers in this work are thin sheets of objects attached to other surfaces, they are almost invisible in the point clouds. This is infeasible if we adopt LiDARTags [18] or calibration boards [15, 16, 17] to provide fiducials. These scenes are challenging due to low overlap, and the overlap rate [2] of each scene is presented in Table II. Competitors and Metrics. Competitors include the state-of-the-art (SOTA) learning-based multiview point cloud registration method, SGHR [1], and the SOTA pairwise geometry-based method, Teaser++ [3]. We employ the root-mean-square errors (RMSEs) [33] as the metric. We also compare the inference time of the methods with an AMD Ryzen 7 5800X CPU. Results and Analysis. The qualitative and quantitative results are presented in Fig. 8 and Table II, respectively. As seen, the proposed method yields the best accuracy and efficiency. SGHR [1] successfully registers the point clouds in scenes 1 to 3 while failing in scenes 4 to 10. Its efficiency is poor on the CPU, while the inference time is around 10s to 20s on an RTX 3090 GPU. Considering that Teaser++ [3] is a pairwise method, we manually select pairs of point clouds with overlap as the inputs (the same scheme adopted in the following tests), yet it still fails in all scenes. The failure cases show SGHR’s limited generalization to unseen scenarios, and Teaser++ struggles with low overlap cases.
Figure 8: Visual comparison with SOTAs (SGHR [1], Teaser++ [3]) on ten scenes. The scenes include the office (1-3), the meeting room (4), the lounge (5,6), the kitchen (7), the office building (8,9), and the thicket (10). Each scene consists of three scans, except scenes 2 and 9, which are composed of two scans.
TABLE II: quantitative comparison with SGHR [1] and Teaser++ [3] on our dataset
Scene ##\## 1 2 3 4 5 6 7 8 9 10
@@ V-B Evaluation of Point Cloud Registration Accuracy
Time (s) 313.1 298.3 336.4 322.3 388.9 419.8 441.5 397.1 327.2 401.2
SGHR[1] $\mathrm{RMSE}_{R}$ (rad) 0.152 0.060 1.198 1.825 1.311 1.696 4.42 2.792 3.034 2.810
@@
Time (s) 846.1 785.0 886.5 825.3 851.7 979.3 983.6 909.9 794.6 950.4
@@
TABLE III: quantitative evaluation of the enhancement of SGHR [1] due to the addition of Livox-3DMatch to the training data.

### v3

@@ IV-C First-level Graph
The objective is to infer the relative poses with the highest possible accuracy using only the necessary low-dimensional information through a simple process. Thus, we construct the first-level graph, shown in Fig. 4, as a weighted graph. The construction process is as follows. When processing the point clouds with the proposed adaptive threshold method one by one, if a marker is detected in a point cloud, we add the corresponding scan node and marker node to the graph along with a weighted edge. The edge weight is the pose estimation point-to-point error ${e}_{pp}$:
@@
Eq. (11) indicates the substitution of the result given by SVD back into the right side of Eq. (8). ${e}_{pp}\in\mathbb{R}^{+}$ is employed as the metric to evaluate the quality of pose estimation of the marker in the corresponding point cloud. We define the first scan as the anchor scan. The local coordinate system of the anchor scan is set as the global coordinate system. Namely, $\{F\}_{1}=\{G\}$. We only consider the relative poses between the anchor scan and each non-anchor scan. Although there may be multiple paths from the given scan to the anchor scan (see Fig. 4), we have already saved the information on pose estimation quality by constructing the first-level graph as a weighted graph. Therefore, we adopt Dijkstra’s algorithm [41] to obtain the shortest path. Then, we compute the relative pose along the shortest path iteratively using Eq. (10). Given that the relative pose computation is achieved with the lowest accumulation of $e_{pp}$, the estimation result achieves the highest possible accuracy. Moreover, the search for the optimal path to propagate poses is based merely on the one-dimensional $e_{pp}$, without incorporating any 6-DOF poses or 3D locations.
@@
Figure 4: An illustration of the first-level graph. After applying the proposed adaptive marker detection to all scans, an exhaustive weighted graph is constructed, with the scans and markers as nodes and the point-to-point errors as edge weights. The aim is to derive the relative pose between each non-anchor scan and the anchor scan with optimal accuracy. However, for a given non-anchor scan, such as $f_{2}$ in this simple case, there may be multiple possible paths in the exhaustive graph leading to the anchor scan ($f_{1}$). Thus, Dijkstra’s algorithm [41] is employed to find the optimal path with the minimum accumulation of point-to-point errors (weights).
IV-D Second-level Graph
To address the MAP problem introduced in Eq. (9), we construct the second-level graph as a factor graph to globally optimize the variables using the initial values obtained from the first-level graph. Specifically, we create the second-level graph in three stages.
Stage One. Suppose that the $j$-th marker is detected in the $i$-th scan, the following six types of nodes are added to the second graph. The variable nodes include:
  • (1)
    Node $\mathbf{T}^{j}$, which refers to the 6-DOF pose of the $j$-th marker w.r.t. $\{G\}$;
  • (2)
    Node $\mathbf{T}_{i}$, which refers to the 6-DOF pose of the $i$-th scan w.r.t. $\{G\}$;
  • (3)
    Nodes $\{\mathbf{p}^{j,1},\mathbf{p}^{j,2},\mathbf{p}^{j,3},\mathbf{p}^{j,4}\}$, which refer to 3D coordinates of the corners of the $j$-th marker w.r.t. $\{G\}$.
The factor nodes include:
  • (4)
    Node $\mathbf{T}^{j}_{i}$, which refers to the measurement of the 6-DOF pose of the $j$-th marker w.r.t. $\{F\}_{i}$;
  • (5)
    Nodes $\{^{j}\mathbf{p}^{j,1},\ ^{j}\mathbf{p}^{j,2},\ ^{j}\mathbf{p}^{j,3},\ ^{j}\mathbf{p}^{j,4}\}$, which refer to the measurement of the 3D coordinates of the corners of the $j$-th marker w.r.t. $\{M\}_{j}$;
  • (6)
    Nodes $\{_{i}\mathbf{p}^{j,1},\ _{i}\mathbf{p}^{j,2},$ ${}_{i}\mathbf{p}^{j,3},$ ${}_{i}\mathbf{p}^{j,4}\}$, which refer to 3D coordinates of the corners of the $j$-th marker w.r.t. $\{F\}_{i}$.
By adding variables representing the 3D coordinates of the corners, we can also optimize the fiducial localization results. The first stage in Fig. 5 shows the added nodes and edges when a marker is detected in a scan.
Stage Two. Thereafter, we traverse all the marker detection results and conduct the operation from Stage One for each detected marker. After processing all the detected markers, the second-level graph becomes the one shown in Stage Two of Fig. 5. Since the operation in this stage essentially consists of repetitions of Stage One, the node types are no different from those in the previous stage.
Stage Three. Considering that the local coordinate system of the first scan, $\{F\}_{1}$, is set as the global coordinate system, $\{G\}$, we add a prior factor that connects to the first (anchor) scan node, $\mathbf{T}^{1}$. Finally, we add factor nodes representing the relative poses between the anchor scan and each non-anchor scan. Up to this point, the factor graph is completed, as shown in Stage Three of Fig. 5.
Following [42, 43], since the factor graph is determined, the posterior probability $P(\Theta\mid\mathcal{Z})$ is factorized as:
@@ IV-D Second-level Graph
where $P^{(k)}$ are the factors in the second graph. We follow the standard pipeline [42] to model them as Gaussians:
@@
where $h_{k}\left(\Theta\right)$ is the measurement function and $z_{k}$ is a measurement. $\|e\|_{\Sigma}^{2}\triangleq e^{T}\Sigma^{-1}e$ denotes the squared Mahalanobis distance with $\Sigma$ being the covariance matrix. Following [28], if $z_{k}\in\mathbb{R}^{3\times 1}$ is a 3D position, $\ominus$ refers to straight subtraction for elements. If $z_{k}\in SE(3)$ is a 6-DOF pose, $\ominus$ generates 6-dimensional Lie algebra (refer to Section III-B) coordinates:
@@
where for a $\mathbf{T}\in SE(3)$, $\mathrm{Rot}(\mathbf{T})$ denotes the rotation matrix $\mathbf{R}\in SO(3)$ and $\mathrm{Trans}(\mathbf{T})$ denotes translation vector $\mathbf{t\in\mathbb{R}^{3\times 1}}$. $\log(\cdot)$ represents the matrix logarithm. $\vee$ is the vee map operator. The detailed introduction of $\vee$ is provided in Section III-B. With the Gaussian modeling, the objective function in Eq. (12) is transformed into a least square problem by applying the negative logarithm:
@@
We utilize the Levenberg-Marquardt algorithm [43] to solve this problem. The acquisition of initial values is introduced in Section IV-C. The noise covariance matrices are determined by the quantitative experiments conducted in [21]. The code implementation, based on Python and C++, is available at https://github.com/yorklyb/L-PR.
Figure 5: The procedures for formulating the second-level graph. The variable nodes are represented by circles and the factor nodes are represented by squares. In Stage One, when a marker is detected in a scan, six types of nodes are added to the graph, including (1) scan pose in $\{G\}$, (2) marker pose in $\{G\}$, (3) corner positions in $\{G\}$, (4) marker pose from $\{M\}$ to $\{G\}$, (5) corner positions in $\{F\}$, and (6) corner positions in $\{M\}$, along with their corresponding edges. In Stage Two, all the marker detection results are traversed, and the operation from Stage One is repeated for each detected marker. In Stage Three, a prior factor connecting the anchor scan is added, along with factors representing the relative poses between the anchor scan and non-anchor scans.
V Livox-3DMatch Dataset
A common approach [1, 2] to evaluating a learning-based point cloud registration model is to train it on 3DMatch [6] and test it on various benchmarks, including 3DMatch [6], ETH [23], and ScanNet [24]. However, the 3DMatch benchmark is mainly constructed from RGB-D camera captures of indoor scenes [6]. In this work, we collect a new dataset named Livox-3DMatch to enrich the training data for learning-based methods. The enrichment contains two key components. Firstly, the sensor we adopt is a Livox MID-40 LiDAR, which has different sampling patterns compared to the RGB-D camera (see Fig. 6(a) and (b)). Thus, it adds point clouds with new features to the training data. Secondly, we selectively sampled some scenes that are absent or rare in 3DMatch, thereby enriching the scenes in the training data. For example, the valid depth range of an RGB-D camera is usually less than ten meters, making it unsuitable for sampling outdoor scenes. In contrast, we collect some outdoor scenes (see Fig. 6(c)) for Livox-3DMatch, considering that a LiDAR can sample objects a few hundred meters away. Moreover, we collect some challenging cases (see Fig. 6(d)) where the overlapping regions are mainly planes, which are rare in 3DMatch [6]. A more detailed introduction to the scenes in Livox-3DMatch is given in Sections VI-B and VI-C.
In Section VI-D, we demonstrate that the proposed Livox-3DMatch can boost the performance of the SOTA learning-based methods [9, 1] on various benchmarks [6, 23, 24].
@@ V Livox-3DMatch Dataset
VI Experimental Results
In this section, we report the experimental results of eight tests to demonstrate the effectiveness and applications of the proposed method from different perspectives.
VI-A Evaluation of the Adaptive Marker Detection Method
In this test, we demonstrate the effectiveness and necessity of the proposed adaptive marker detection method.
Data. We place the Livox MID-40 LiDAR and 69.2 cm $\times$ 69.2 cm ArUco marker(s) in three different scenes, as shown in Fig. 7. The three scenes are: between groups of buildings, in open outdoor areas, and in large indoor parking lots. In each scene, we collect point clouds and test the proposed adaptive threshold method at different relative positions.
Results and Analysis. Table I presents the results, where x and y indicate the relative position of the LiDAR in the marker (ID 4) coordinate system. Specifically, in scene Fig. 7(c), there are two markers (ID 4 and ID 1), so the column for $\lambda^{*}$ contains two values. As seen in the table, $\lambda^{*}$ varies significantly as the relative position changes, demonstrating the necessity of the proposed adaptive marker detection method. In addition, the results shown in Fig. 7(c) also illustrate the necessity of the memory queue design, considering that a single $\lambda^{*}$ might not be applicable to all markers in the same scene. Please note that the intention of reporting the values of the optimal threshold in Table I is to demonstrate that a constant threshold, as adopted in [21], is not applicable to this task. However, the major concern of the adaptive threshold marker detection algorithm is to detect all markers in a scene. Namely, the marker detection results saved in the memory queue are the most important output rather than the thresholds.
@@ VI-A Evaluation of the Adaptive Marker Detection Method
TABLE I: Demonstration of the necessity of the proposed adaptive marker detection algorithm.
@@
VI-B Evaluation of Point Cloud Registration Accuracy
In this test, we evaluate the proposed method against competitors in terms of point cloud registration accuracy.
Data. Given that the existing point cloud registration benchmarks [6, 23, 24] lack fiducial markers in the scenes, we construct a new dataset, as shown in Fig. 8, with the Livox MID-40. As illustrated in the caption of Fig. 8, our dataset covers various indoor and outdoor scenes. Indoors, multiple 16.4 cm × 16.4 cm AprilTags [12] are positioned in the environment. Outdoors, multiple 69.2 cm × 69.2 cm ArUcos [13] are placed in the environment. Note that since the LiDAR fiducial markers in this work are thin sheets of objects attached to other surfaces, they are invisible in the 3D point clouds. This is infeasible if we adopt LiDARTags [20] or calibration boards [17, 18, 19] to provide fiducials. These scenes are challenging due to low overlap, and the overlap rate [2] of each scene is presented in Table II. The ground truth poses are obtained by manually registering through CloudCompare [44], a popular 3D point cloud processing software.
Competitors and Metrics. Competitors include the latest state-of-the-art (SOTA) multiview point cloud registration methods, MDGD [9] (RA-L24), SGHR [1] (CVPR23), and the SOTA pairwise methods, SE3ET [7] (RA-L24), GeoTrans [8] (TPAMI23), and Teaser++ [3] (T-RO20). All the competitors are learning-based methods, except for Teaser++, which is a geometry-based method. For all pairwise methods, we manually select pairs of point clouds that have overlap as the inputs. We employ the root-mean-square errors (RMSEs) [45] as the metric. We also compare the inference time of the methods with an AMD Ryzen 7 5800X CPU.
Figure 8: A comparison with SOTA methods, including MDGD [9], SGHR [1], SE3ET [7], GeoTrans [8], and Teaser++ [3], on ten scenes. The scenes include the office (1-3), the meeting room (4), the lounge (5,6), the kitchen (7), the office building (8,9), and the thicket (10). Each scene consists of three scans, except scenes 2 and 9, which are composed of two scans.
TABLE II: quantitative comparison with MDGD [9], SGHR [1], SE3ET [7], GeoTrans [8], and Teaser++ [3] on our dataset.
Scene ##\## 1 (Office) 2 (Office) 3 (Office) 4 (Meeting Room) 5 (Lounge) 6 (Lounge) 7 (Kitchen) 8 (Building) 9 (Building) 10 (Thicket)
@@ VI-B Evaluation of Point Cloud Registration Accuracy
(T-RO20) Time (s) 313.1 298.3 336.4 322.3 388.9 419.8 441.5 397.1 327.2 401.2
GeoTrans [8] $\mathrm{RMSE}_{R}$ (rad) 1.331 1.451 1.935 1.632 1.532 1.471 1.198 1.355 1.377 1.526
$\mathrm{RMSE}_{T}$ (m) 1.502 1.671 1.912 1.552 1.325 1.235 1.132 1.730 1.554 1.531
(TPAMI23) Time (s) 91.8 87.3 95.1 93.9 105.3 122.3 127.3 110.5 91.1 121.4
SE3ET [7] $\mathrm{RMSE}_{R}$ (rad) 0.925 0.834 1.776 1.358 1.381 1.352 1.077 1.156 1.232 1.376
$\mathrm{RMSE}_{T}$ (m) 0.996 1.113 1.736 1.273 1.207 1.130 1.095 1.469 1.413 1.392
(RA-L24) Time (s) 116.8 94.6 99.5 95.7 115.7 136.0 139.3 123.5 97.3 132.0
SGHR [1] $\mathrm{RMSE}_{R}$ (rad) 0.152 0.060 1.198 1.825 1.311 1.696 3.42 2.792 3.034 2.810
@@
(CVPR23) Time (s) 846.1 785.0 886.5 825.3 851.7 979.3 983.6 909.9 794.6 950.4
MDGD [9] $\mathrm{RMSE}_{R}$ (rad) 0.032 0.062 1.457 0.933 0.679 1.715 0.856 0.722 0.651 0.779
$\mathrm{RMSE}_{T}$ (m) 0.015 0.009 0.352 1.035 0.113 0.422 0.337 1.553 0.686 0.939
(RA-L24) Time (s) 626.1 573.3 645.5 619.2 612.0 726.4 730.5 689.6 590.6 717.4
@@
Results and Analysis. The qualitative and quantitative results are presented in Fig. 8 and Table II, respectively. The pairwise methods [7, 8, 3], although provided manually selected pairs of point clouds with overlap, struggle in all scenes. This is because they are tailored for extracting geometric features either in a learning-based manner [7, 8] or through conventional geometry [3]. Thus, when the overlap ratio between point clouds is low and the overlapped regions lack sufficient geometric features, these methods struggle to find features and utilize them to align point clouds. In contrast, the multiview point cloud registration methods [9, 1] show some successful cases. In particular, SGHR [1] successfully registers the point clouds in scenes 1 to 3 and one point cloud pair in scene 4. MDGD [9], built upon the framework of SGHR [1] but developing a new matching distance-based overlap estimation module, aligns scenes 1 and 2 with decent accuracy and also successfully registers one pair of point clouds in scenes 4, 5, 7, and 10. The reason is that, in addition to learning pairwise registration, the multiview point cloud registration methods [9, 1] also learn to further optimize or refine the pose graph constructed using pairwise registration results. Despite this, they fail in other scenes. Reviewing their success cases, it is found that these indoor scenes are similar to those in their training dataset [6]. While failure cases, such as scenes 4 to 10, are rare or lacking in the training dataset. This indicates that their generalization to unseen scenarios is limited due to the out-of-distribution problem. The proposed method achieves the best performance. Specifically, our method does not require the overlapping regions to have explicit geometric features, thanks to the thin-sheet format of our LiDAR fiducial markers. Moreover, by virtue of the proposed adaptive marker detection algorithm, our method can robustly align any novel indoor or outdoor scenes with thin-sheet markers. The proposed method also yields the best efficiency as it focuses on registering point clouds through LiDAR fiducial markers rather than analyzing the entire point clouds. The process of our method takes several dozen seconds, while other methods need several minutes on the AMD Ryzen 7 5800X CPU.
Figure 9: An illustration of the experimental setup and a visual comparison of the proposed method against the SOTA methods (MDGD [9], SGHR [1], SE3ET [7], GeoTrans [8], and Teaser++ [3]) regarding instance reconstruction.
Figure 10: A visual comparison of the ground truth shape (top) and the reconstructed shape using the proposed method (bottom).
VI-C Application 1: 3D Asset Collection from Sparse Scans
Collecting the complete shape of a novel object [25] from sparse observations is advantageous, as sparse scans require less labor and offer better efficiency. However, it is also a challenging task due to the low overlap between scans. In this test, we evaluate the instance reconstruction quality of the proposed method.
Data. The experimental setup is depicted in Fig. 9. Four 69.2 cm $\times$ 69.2 cm ArUco [13] markers are placed in the environment. We employ the Livox MID-40 LiDAR to scan the vehicle from five significantly different viewpoints. As shown in Fig. 9, this is also a challenging low overlap case, with the average and minimum overlap rates being 13.39% and 1.02%, respectively. To evaluate the instance reconstruction quality, we use Supervisely [46], a popular 3D annotation tool, to extract the point cloud of the vehicle from the reconstruction result. Since the vehicle’s make and model (Mercedes-Benz GLB) are known, a high-fidelity 3D model from a 3D assets website serves as the ground truth shape.
Competitors and Metrics. The competitors are the SOTA point cloud registration methods, including MDGD [9], SGHR [1], SE3ET [7], GeoTrans [8], and Teaser++ [3]. Pairs of point clouds with overlapping regions are manually provided to the pairwise methods [7, 8, 3]. MDGD [9], SGHR [1], and our method directly process the set of point clouds. In terms of metrics, we follow [25] and employ Chamfer Distance and Recall between the ground truth shape and the reconstructed shape to evaluate reconstruction quality.
TABLE III: quantitative evaluation of the enhancement of SGHR [1] and MDGD [9] due to the addition of Livox-3DMatch to the training.